Grade 11General Physics

Q.A boat is able to move through still water at 20m/s.It makes a round trip to a town 3.0km upstream.If the river flows at 5m/s,the time required for this round trip is:
A)120s
B)150s
C)200s
D)320s

jiten
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
Dear Jiten
 
V = d/t 

so t = d/V 

On the upstream part the boat is only going 20 m/s - 5 m/s = 15 m/s because the river pushing back. 

t = 3000 m / 15 m/s = 200 s 

And on the strip back the speed is 20 m/s + 5 m/s = 25 m/s because the river is now pushing you from behind adding to the speed 

t = 3000 n / 25 m/s = 120 s 

total time = 200 s + 120 s = 320 sec 
 
Regards
Arun (askIITians forum expert)